Solve for X. The measure of angle AOB = 4x-2. The measure of angle BOC = 5x+10. The measure of angle COD = 2x+14.
I thought that if I added the 3 angles together (11X+22) that the solution would equal angle AOD. However, I can't seem to solve for x when I do that. What am I doing wrong? thanks.

Answer by venugopalramana(3286)

Solve for X. The measure of angle AOB = 4x-2. The measure of angle BOC = 5x+10. The measure of angle COD = 2x+14.
I thought that if I added the 3 angles together (11X+22) that the solution would equal angle AOD. However, I can't seem to solve for x when I do that. What am I doing wrong? thanks.
YOU HAVE NOT GIVEN ANY FIGURE NOR EXPLAINED WHAT/WHERE ARE A,B,C,D,O..ASSUMING AS GIVEN BELOW ,I AMM TRYING TO EXPLAIN.BUT PLEASE GIVE COMPLETE PROBLEM TO GET CORRECT FEED BACK
YOU DID WELL BUT MISSED A POINT IN THE END.WE ARE TALKING OF A QUADRILATERAL ABCD,WITH O AS A POINT ITS MIDDLE,JOINING OA,OB,OC AND OD AND FINDING THE ANGLE AOD....TO SOLVE YOU SHOULD HAVE AN EQUATION..YOU SUMMED UP THE 3 ANGLES AND GOT 11X+22..OK...BUT IT IS NOT ANGLE AOD .THAT PLUS ANGLE AOD COMPLETE A CIRCULAR ANGLE AROUND O..THE KEY FACT IS THAT IT SHOULD EQUAL 360 WHICH GIVE YOU THE EQN. TO SOLVE ..SO
11X+22+AOD=360
ANGLE AOD=360-11X-22=338-11X